# Npix2Cpix: A GAN-based Image-to-Image Translation Network with Retrieval-Classification Integration for Watermark Retrieval from Historical Document Images

### v1

@@ Abstract
The identification and restoration of ancient watermarks have long been a major topic in codicology and history. Classifying historical documents based on watermarks can be difficult due to the diversity of watermarks, crowded and noisy samples, multiple modes of representation, and minor distinctions between classes and intra-class changes. This paper proposes a U-net-based conditional generative adversarial network (GAN) to translate noisy raw historical watermarked images into clean, handwriting-free images with just watermarks. Considering its ability to perform image translation from degraded (noisy) pixels to clean pixels, the proposed network is termed as Npix2Cpix. Instead of employing directly degraded watermarked images, the proposed network uses image-to-image translation using adversarial learning to create clutter and handwriting-free images for restoring and categorizing the watermarks for the first time. In order to learn the mapping from input noisy image to output clean image, the generator and discriminator of the proposed U-net-based GAN are trained using two separate loss functions, each of which is based on the distance between images. After using the proposed GAN to pre-process noisy watermarked images, Siamese-based one-shot learning is used to classify watermarks. According to experimental results on a large-scale historical watermark dataset, extracting watermarks from tainted images can result in high one-shot classification accuracy. The qualitative and quantitative evaluation of the retrieved watermarks illustrates the effectiveness of the proposed approach.
Index Terms:
Historical Watermark, Watermark retrieval, Generative Adversarial Network, Siamese NetworkI Introduction
@@
Watermark recognition from historical images and retrieval is a crucial task for historians and archivists in order to classify and explore the origin of a historical document. Ancient watermarks help the users to estimate the dates of creation of the watermarked documents (texts or drawings) without any explicit chronological indication [1] [2], as well as to determine the probable location of the document’s origin [3]. For archivists, historians, curators, and anyone else interested in the trade of written documents, being able to recognize the watermark on historical papers and obtain the clean version of that watermark, makes it possible to precisely date the paper and determine the paper manufacturer by looking through existing watermark design catalogs (booksellers, antiquarians, etc.). Due to natural deterioration as well as man-made catastrophes, there is a danger that the information contained in the old documents would be lost after many years.
@@
Even if the documents are preserved carefully, there is a great possibility that they will continue to deteriorate over time. Watermarks are essential for gathering crucial data, classifying, and categorizing documents. That is why, in recent years, it has become more essential than ever to be able to accurately identify historical watermarks in documents. Even though this is very useful, automatic watermark detection is still very difficult that has not yet been fully resolved. In addition to being a useful resource in the preservation of cultural heritage collections, historical watermarks serve as a source of information for archivists and historians. Due to the variety of patterns used in these watermarks, they are very difficult to identify and recognize. The identification issue is further complicated by the presence of handwriting on the watermarks and various types of noises [4]. Hand tracing and backlit photography are two traditional methods of watermark retrieval [5]. A number of research works have emphasized localizing and extracting a watermark pattern from a back-lit image, rarely incorporating aligned reflected light images [6], [7], [8]. Aside from analog watermarks on historical documents, digital watermarking became popular in the late 1990s in order to embed robust digital marks into diverse materials and images to identify the origin, owner, rights, and integrity, as well as to ensure security. In addition to steganography, spread spectrum communications technology, and noise theory-based generic watermarking systems [9], there has been a lot of progress in computer graphics-based robust watermarking tools, such as block-based [10] and triangle-based ones [11]. Automatic image retrieval has long been recognized as an intriguing method for historical document study. In the late 1990s, Rauber et al. [12] and Belov et al. [13] discussed watermark retrieval as a use case for content-based image retrieval (CBIR) approaches. A CBIR system originally developed for trademark retrieval has been applied to tracings and radiographies of historical watermarks in the Northumbria watermarks archive project [14]. Compared to radiographs, the system performed significantly better on tracings [15]. Two aspects associated with document enhancement are highlighted in [16], namely the recovery of damaged documents and watermark removal. In [17], many of the existing documents are digitized using smartphone cameras. These are highly vulnerable to capturing distortions (perspective angle, shadow, blur, warping, etc.), making them hard to read by a human or by an OCR engine. Lucene Image REtrieval (LIRE) is a Java open-source program that was first made available in 2006 [18]. Based on the Lucene search engine, in the LIRE program, both global and local features are used. Low-level color, texture, and certain mixed characteristics make up the global features. Image retrieval performance is demonstrated by using the global features of LIRE’s free demo edition (Fuzzy color, texture histogram, and MPEG7 color layout). Recently in [19], considering as an application of the proposed method, digital watermarks from images have been eliminated. Initially, watermark-related nodes are chosen using the various schemes proposed here. Finally, local manipulation is employed to delete all associated control points with the watermark. In [20], a blind color picture watermarking system with good spatial performance is proposed to achieve effective copyright protection of color images by combining the advantages of spatial-domain and frequency-domain watermarking schemes. Image watermarking algorithms such as 1D empirical mode decomposition (EMD) and dimensional reduction via Hilbert curve [21], watermarking system with a memristor-based hyperchaotic oscillator [22], two-stage fuzzy inference system [23], and others have recently gained popularity for the purpose of protecting copyright and preventing information from being used without authorization. Deep learning algorithms are getting more attention for historical document analysis. Identifying an abstract representation—such as drawings, paintings, plush toys, sculptures, signs, or silhouettes—is an easy and frequent task for humans, but computers continue to struggle with it. Histogram-based descriptors [24] have been used in older studies like [25], whereas the use of machine learning techniques like dictionary learning [26] and neural networks [27] are noticeable in more recent works. In [26], a retrieval strategy for categorizing watermarks is proposed based on local features and a dictionary that is trained on a large amount of data. The quest for comparable deep learning-based historical watermark identification and recognition systems has been made by several researchers. The technique provided in [27], solves the problem of identifying the same object even when it is depicted in a variety of manners also known as the cross-depiction issue, on a historical watermark dataset [28]. This dataset is not publicly available and contains 106,000 watermark pictures.
@@
Recently, a research on image retrieval for watermarks is jointly conducted by the Institut de Recherche et d’Histoire des Textes (IRHT), the Institut National de Recherche en Informatique et en Automatique (INRIA), and the Ecole Nationale des Chartes and results are reported in [35]. A large public dataset with over 6000 original images is published here, allowing researchers to tackle scenarios of practical importance at scale. A matching score and feature fine-tuning technique based on filtering local matches is also introduced here using spatial consistency. In [36], a new publicly accessible online tool for automated watermark identification is proposed based on [35]. In [37], a spatially consistent feature matching and learning algorithm is proposed for historical image analysis. Here, an image similarity score based on geometric validation of mid-level features is defined. In this work, it is also demonstrated how weak or no supervision may be used to fine-tune out-of-the-box features for the target dataset using spatial consistency. Italian papers of the late Middle Ages with watermarks are classified using machine learning in [2]. Here, a new spatial-temporal methodology for tracing regional paper mill production and regional recipes is also introduced. Another issue with watermark detection or classification is the multi-domain representations of data, that can be solved using recent studies like [38]. Recently, in [39], a Siamese-based one-shot network for automatic watermark recognition from the historical image dataset has been proposed. Here, a thresholding-based image enhancement technique is employed, that strongly depends on the chosen threshold value. The introduction of the one-shot classification in historical watermark classification, proposed in this paper, is very useful for the low data regime, which has great demand.
@@
The common problem in watermark detection is the scarcity of clean examples to test newer watermarks. There are also a lot of messy and distorted samples. These watermarks can be challenging to detect at times because of the multiple modes of representation and both small discrepancies between classes and substantial intra-class variations. As a solution to such problems, in this paper, a generative adversarial network (GAN), named as Npix2Cpix, is proposed to translate raw noisy images to their cleaned counterpart. Although deep learning-based models have been used previously to remove entities from images [40], [41] image-to-image translation applications to extract a specific piece of information or pattern from historical document images have not yet been explored. The phrase ”translating” refers to converting an input noisy image into a corresponding clean output image, that can be used to describe a variety of issues in image processing, computer graphics, and computer vision. A U-net-based architecture serves as the generator of the GAN, that has been proposed here, reconstructing the clear image from the noisy one. Both the generated image and the noisy image are compared with the instance’s ground truth. The generator is trained to produce images that are somewhat similar to the ground truth, while the discriminator is taught to discern between ground truth and generated images. To update the parameters of the network, a combined loss function is employed. For the categorization of historical watermarks, a Siamese-based one-shot classification pipeline is proposed. Here, two identical neural networks are used to process both support and query images, using the same structure, weights, and parameters. Cosine similarity is used in this case to calculate similarity.
The main contributions of this paper are:
@@
A U-net-based generative adversarial image-to-image translation network for watermark retrieval from historical documents.
@@
The proposed U-Net-based conditional GAN is not simply used to generate new synthetic instances of images, rather it focuses on enhancing the clarity and quality of watermarked historical document images, facilitating subsequent watermark classification and analysis. Hence the generator network of the proposed U-Net-based GAN is specially designed to map a noisy watermarked image to a generated noise-reduced watermarked image. In particular, different modifications are incorporated in the U-Net structure used in the generator network, such as an increased number of up-sampling and down-sampling blocks, a dropout layer in the up-sampling path, and the Tanh activation function in the final layer.
In the proposed Siamese-based one-shot classification network, in order to handle the extremely degraded and might not be fully denoised historical documents, unlike traditional Siamese networks, different pretrained networks, such as ResNet and EfficientNet, are integrated as the feature extractors along with cosine distance-based similarity measurement. By combining the modified U-Net-based conditional GAN for noisy image to noise-reduced image translation and a subsequent one-shot classification scheme using the proposed Siamese network, the proposed method effectively addresses the complex challenges of classifying ancient watermarked document images.
@@ II Dataset Description
III Proposed Method
Traditional methods for watermark retrieval and classification usually input the noisy image directly into the classifier network, which then does classification, and recent works for denoising documents and classification are based on deep residual networks and auto-encoders [42], [43]. As a result, additional effort is necessary to expend into developing a robust classification network. This is because, in this case, the classification network will have to deal with both the classification and noise reduction problems. Thus, the model could not provide sufficient performance, especially if the input image is severely degraded. Other techniques focus specially on the watermarked regions for retrieval. As ancient documents with watermarks contain handwriting and additional noise, however, there is a possibility that it ends up enhancing the noise and writing of that region as well. On the other hand, a crucial part of this research is to handle one-shot classification where there is less scope for getting a huge amount of data for traditional classification. As a solution to the problems associated with traditional methods, this research introduces two major schemes. Initially, in an effort to mitigate the impact of noise and handwritten artifacts, a U-net-based conditional GAN is proposed in this paper. This network facilitates the transformation of noisy historical images into their clean counterparts by effectively eliminating handwriting and extraneous noise, thereby constituting an image-to-image translation scheme. In order to accomplish noise removal and watermark extraction, the suggested image-to-image translation approach with U-net-based GAN has been named as Npix2Cpix. Furthermore, to address the limitation imposed by the scarcity of training data, a modified Siamese-based one-shot classification network is used. A GAN-based network employed prior to the one-shot classification stage will be very efficient, as instead of a noisy image, a noise-reduced as well as writing removed image needs to be handled in the classifier. The overall idea proposed in this paper can be summarized into two phases, they are - Noise and writing removal using U-net-based GAN and One-shot classification. The problem statement along with suggested noise and handwriting removal with classification pipelines are described in the following subsections.
III-A Problem Statement
Historical images, to be more specific, historical watermark document images usually have some additional challenges compared to traditional computer vision tasks. Specifically, there are three major technical challenges in the watermark retrieval from historical document images, namely-
@@ III-A Problem Statement
The estimation of background within historical document images holds significant importance due to the presence of watermarks typically situated in the background section of the image. However, detecting the background in such instances represents a notably challenging task. The background of historical images is often noisy, blurry, of different brightness levels, spotted, and non-homogeneous, as can be seen in Fig. 1. Here, noise refers to the discontinuity and non-uniformity of the background paper. The degradation of watermarked documents with time, often variations in different locations of the same image as well as images from different sources, affects the brightness, clarity, sharpness, and overall colors of the watermark background. These characteristics make it difficult for the model to estimate the historical image’s background. The watermarks to be retrieved from ancient document images vary significantly in terms of their types, patterns, and level of noise corruption. This watermarks can occasionally be extremely faded, distorted, and discontinuous due to the aging or data acquisition process. Besides, the patterns of the watermarks have a huge variation in their size, complexity, location, design, thickness, and clarity. However, observing images in Fig.1, it can be agreed that the biggest challenge to extracting significant features from these images is the blackish-brown handwriting, which is randomly spread all over most of the images. To be more specific, the handwriting on the top of the watermarks creates additional complexity in retrieving watermarks. Besides, there is a huge variation in complexity, thickness, color, and depth (Dark black, Light brown, etc.), size, and style of the letters. And the biggest challenge for the model will be to generalize this variation. Besides, as the images are obtained from different sources, the handwriting not only varies in writing style and typographical characteristics, they might be of such fonts that do not exist at present and in no way related from image to image. These three major issues have to be handled in this research work.
III-B Image to Image Translation: Npix2Cpix

### v3

@@ Abstract
The identification and restoration of ancient watermarks have long been a major topic in codicology and history. Classifying historical documents based on watermarks is challenging due to their diversity, noisy samples, multiple representation modes, and minor distinctions between classes and intra-class variations. This paper proposes a modified U-net-based conditional generative adversarial network (GAN) named Npix2Cpix to translate noisy raw historical watermarked images into clean, handwriting-free watermarked images by performing image translation from degraded (noisy) pixels to clean pixels. Using image-to-image translation and adversarial learning, the network creates clutter-free images for watermark restoration and categorization. The generator and discriminator of the proposed GAN are trained using two separate loss functions, each based on the distance between images, to learn the mapping from the input noisy image to the output clean image. After using the proposed GAN to pre-process noisy watermarked images, Siamese-based one-shot learning is employed for watermark classification. Experimental results on a large-scale historical watermark dataset demonstrate that cleaning the noisy watermarked images can help to achieve high one-shot classification accuracy. The qualitative and quantitative evaluation of the retrieved watermarked image highlights the effectiveness of the proposed approach.
keywords:
Historical Watermark , Watermark Retrieval , Generative Adversarial Network , Siamese Network[inst1]organization=Department of Electrical and Electronic Engineering,addressline=BUET, city=Dhaka, postcode=1000, country=Bangladesh
[…]
1 Introduction
@@
Watermark recognition from historical images and retrieval is a crucial task for historians and archivists in order to classify and explore the origin of a historical document. Ancient watermarks help the users to estimate the dates of creation of the watermarked documents (texts or drawings) without any explicit chronological indication [1] [2], as well as to determine the probable location of the document’s origin [3]. For archivists, historians, curators, and anyone else interested in the trade of written documents, being able to recognize the watermark on historical papers and obtain the clean version of that watermark, makes it possible to precisely date the paper and determine the paper manufacturer by looking through existing watermark design catalogs (booksellers, antiquarians, etc.). Due to natural deterioration as well as man-made catastrophes, there is a danger that the information contained in the old documents would be lost after many years. Even if the documents are preserved carefully, there is a great possibility that they will continue to deteriorate over time. Ancient watermarks are essential for gathering crucial data, and categorizing historical documents. That is why, in recent years, it has become more essential than ever to be able to accurately identify historical watermarks in documents. Despite its great usefulness, automatic watermark detection remains highly challenging and remains unresolved. In addition to being a useful resource in the preservation of cultural heritage collections, historical watermarks serve as a source of information for archivists and historians. Due to the variety of patterns used in these watermarks, they are very difficult to identify and recognize. The identification issue is further complicated by the presence of handwriting on the watermarks and various types of noises [4].
@@
Hand tracing and backlit photography are two traditional methods of watermark retrieval [5]. A number of research works have emphasized localizing and extracting a watermark pattern from a back-lit image, rarely incorporating aligned reflected light images [6], [7], [8]. Aside from analog watermarks on historical documents, digital watermarking became popular in the late 1990s in order to embed robust digital marks into diverse materials and images to identify the origin, owner, rights, and integrity, as well as to ensure security. In addition to steganography, spread spectrum communications technology, and noise theory-based generic watermarking systems [9], there has been a lot of progress in computer graphics-based robust watermarking tools, such as block-based [10] and triangle-based ones [11]. Automatic image retrieval has long been recognized as an intriguing method for historical document study. In the late 1990s, Rauber et al. [12] and Belov et al. [13] discussed watermark retrieval as a use case for content-based image retrieval (CBIR) approaches. A CBIR system originally developed for trademark retrieval has been applied to tracings and radiographies of historical watermarks in the Northumbria watermarks archive project [14]. Compared to radiographs, the system performed significantly better on tracings [15]. Two aspects associated with document enhancement are highlighted in [16], namely the recovery of damaged documents and watermark removal. In [17], many of the existing documents are digitized using smartphone cameras. These are highly vulnerable to capturing distortions (perspective angle, shadow, blur, warping, etc.), making them hard to read by a human or by an OCR engine.
Lucene Image REtrieval (LIRE) is a Java open-source program that was first made available in 2006 [18]. Based on the Lucene search engine, in the LIRE program, both global and local features are used. Low-level color, texture, and certain mixed characteristics make up the global features. Image retrieval performance is demonstrated by using the global features of LIRE’s free demo edition (Fuzzy color, texture histogram, and MPEG7 color layout). Recently in [19], considering as an application of the proposed method, digital watermarks from images have been eliminated. Initially, watermark-related nodes are chosen using the various schemes proposed here. Finally, local manipulation is employed to delete all associated control points with the watermark. In [20], a blind color picture watermarking system with good spatial performance is proposed to achieve effective copyright protection of color images by combining the advantages of spatial-domain and frequency-domain watermarking schemes. Image watermarking algorithms such as 1D empirical mode decomposition (EMD) and dimensional reduction via Hilbert curve [21], watermarking system with a memristor-based hyperchaotic oscillator [22], two-stage fuzzy inference system [23], and others have recently gained popularity for the purpose of protecting copyright and preventing information from being used without authorization.
Deep learning algorithms are getting more attention for historical document analysis. Identifying an abstract representation—such as drawings, paintings, plush toys, sculptures, signs, or silhouettes—is an easy and frequent task for humans, but computers continue to struggle with it. Histogram-based descriptors [24] have been used in older studies like [25], whereas the use of machine learning techniques like dictionary learning [26] and neural networks [27] are noticeable in more recent works. In [26], a retrieval strategy for categorizing watermarks is proposed based on local features and a dictionary that is trained on a large amount of data. The quest for comparable deep learning-based historical watermark identification and recognition systems has been made by several researchers. The technique provided in [27], solves the problem of identifying the same object even when it is depicted in a variety of manners also known as the cross-depiction issue, on a historical watermark dataset [28]. This dataset is not publicly available and contains 106,000 watermark pictures.
@@
Recently, research on image retrieval for watermarks has been jointly conducted by the Institut de Recherche et d’Histoire des Textes (IRHT), the Institut National de Recherche en Informatique et en Automatique (INRIA), and the Ecole Nationale des Chartes and results are reported in [35]. A large public dataset with over 6000 original images is published here, allowing researchers to tackle scenarios of practical importance at scale. A matching score and feature fine-tuning technique based on filtering local matches is also introduced here using spatial consistency. In [36], a new publicly accessible online tool for automated watermark identification is proposed based on [35]. In [37], a spatially consistent feature matching and learning algorithm is proposed for historical image analysis. Here, an image similarity score based on geometric validation of mid-level features is defined. In this work, it is also demonstrated how weak or no supervision may be used to fine-tune out-of-the-box features for the target dataset using spatial consistency. Italian papers of the late Middle Ages with watermarks are classified using machine learning in [2]. Here, a new spatial-temporal methodology for tracing regional paper mill production and regional recipes is also introduced. Another issue with watermark detection or classification is the multi-domain representations of data, that can be solved using recent studies like [38]. Recently, in [39], a Siamese-based one-shot network for automatic watermark recognition from the historical image dataset has been proposed. Here, a thresholding-based image enhancement technique is employed, that strongly depends on the chosen threshold value. The introduction of the one-shot classification in historical watermark classification, proposed in this paper, is very useful for the low data regime, which has great demand.
@@
Reviewing the existing studies it can be agreed that the common problem in watermark detection is the scarcity of clean examples to test newer watermarks. There are also a lot of messy and distorted samples. These watermarks can be challenging to detect at times because of the multiple modes of representation, small discrepancies between classes, and substantial intra-class variations. As a solution to such problems, in this paper, a conditional generative adversarial network (GAN), named Npix2Cpix, is proposed to translate raw, noisy images to their cleaned counterparts. Although deep learning-based models have been used previously to remove entities from images [40], [41] image-to-image translation applications to extract a specific piece of information or pattern from historical document images have not yet been explored. A recent approach improves multimodal image-to-image translation by using a shared encoder and generator for all domains, enhancing feature extraction and performance in complex tasks [42]. Similarly, another method enhances change detection in optical and SAR images by aligning them into a common feature space, achieving higher accuracy and robustness [43].
The phrase ”translating” refers to converting an input noisy image into a corresponding clean output image, that can be used to describe a variety of issues in image processing, computer graphics, and computer vision. A U-net-based architecture serves as the generator of the GAN, that has been proposed here, reconstructing the clear image from the noisy one. Both the generated image and the noisy watermarked image are compared with the instance’s ground truth. The generator is trained to produce images that are somewhat similar to the ground truth, while the discriminator is taught to discern between ground truth and generated images. To update the parameters of the network, a combined loss function is employed. Subsequently, for the categorization of historical watermarks, a Siamese-based one-shot classification pipeline is proposed. Here, two identical neural networks are used to process both support and query images, using the same structure, weights, and parameters. Cosine similarity is used in this case to calculate similarity. The main contributions of this paper are:
1.
@@
A modified U-net-based generative adversarial image-to-image translation network for watermark retrieval from historical documents.
2.
@@
The proposed U-Net-based conditional GAN is not simply used to generate new synthetic instances of images, rather it focuses on enhancing the clarity and quality of watermarked historical document images, facilitating subsequent watermark classification and analysis. Hence the generator network of the proposed U-Net-based GAN is specially designed to map a noisy watermarked image to a generated noise-reduced watermarked image. In particular, different modifications are incorporated in the U-Net structure used in the generator network, such as an increased number of up-sampling and down-sampling blocks, a dropout layer in the up-sampling path, and the Tanh activation function in the final layer. In the proposed Siamese-based one-shot classification network, various pretrained models, including ResNet, MobileNet, and EfficientNet, are integrated as feature extractors, along with cosine distance-based similarity measurements. This approach is designed to handle extremely degraded historical documents that may not be fully denoised, distinguishing it from traditional Siamese networks. By combining the modified U-Net-based conditional GAN for noisy image to noise-reduced image translation and a subsequent one-shot classification scheme using the proposed Siamese network, the proposed method effectively addresses the complex challenges of classifying ancient watermarked document images.
@@ 2 Dataset Description
3 Methodology
Traditional methods for watermark retrieval and classification usually input the noisy image directly into the classifier network, which then does classification, and recent works for denoising documents and classification are based on deep residual networks and auto-encoders [44], [45]. As a result, additional effort is necessary to expend into developing a robust classification network. This is because, in this case, the classification network will have to deal with both the classification and noise reduction problems. Thus, the model could not provide sufficient performance, especially if the input image is severely degraded. Other techniques focus specifically on the watermarked regions for retrieval. As ancient documents with watermarks contain handwriting and additional noise, there is a possibility that it ends up enhancing the noise and writing of that region as well. On the other hand, a crucial part of this research is to handle one-shot classification where there is less scope for getting a huge amount of data for traditional classification. As a solution to the problems associated with traditional methods, this research introduces two major schemes. Initially, in an effort to mitigate the impact of noise and handwritten artifacts, a U-net-based conditional GAN is proposed in this paper. This network facilitates the transformation of noisy historical images into their clean counterparts by effectively eliminating handwriting and extraneous noise, thereby constituting an image-to-image translation scheme. In order to accomplish noise removal and watermark extraction, the suggested image-to-image translation approach with U-net-based GAN has been named Npix2Cpix. Furthermore, to address the limitation imposed by the scarcity of training data, a modified Siamese-based one-shot classification network is used. A GAN-based network employed prior to the one-shot classification stage is very efficient, as instead of a noisy image, a noise-reduced as well as writing removed image needs to be handled in the classifier. The overall idea proposed in this paper can be summarized into two phases, they are - noise and writing removal using U-net-based GAN and one-shot classification. The problem statement along with suggested noise and handwriting removal with classification pipelines are described in the following subsections.
3.1 Problem Statement
Historical images, to be more specific, historical watermarked document images usually have some additional challenges compared to traditional computer vision tasks. Specifically, there are three major technical challenges in the watermark retrieval from historical document images, namely-
1.
@@ 3.1 Problem Statement
The estimation of background within historical document images holds significant importance due to the presence of watermarks typically situated in the background section of the image. However, detecting the background in such instances represents a notably challenging task. The background of historical images is often noisy, blurry, of different brightness levels, spotted, and non-homogeneous, as can be seen in Fig. 1. Here, noise refers to the discontinuity and non-uniformity of the background paper. The degradation of watermarked documents with time, often variations in different locations of the same image as well as images from different sources, affects the brightness, clarity, sharpness, and overall colors of the watermark background. These characteristics make it difficult for the model to estimate the historical image’s background. The watermarks to be retrieved from ancient document images vary significantly in terms of their types, patterns, and level of noise corruption. These watermarks can occasionally be extremely faded, distorted, and discontinuous due to the aging or data acquisition process. Besides, the patterns of the watermarks have a huge variation in their size, complexity, location, design, thickness, and clarity. However, observing images in Fig.1, it can be agreed that the biggest challenge to extracting significant features from these images is the blackish-brown handwriting, which is randomly spread all over most of the images. To be more specific, the handwriting on the top of the watermarks creates additional complexity in retrieving watermarks. Besides, there is a huge variation in complexity, thickness, color, depth (Dark black, Light brown, etc.), size, and style of the letters. And the biggest challenge for the model will be to generalize this variation. Besides, as the images are obtained from different sources, the handwriting not only varies in writing style and typographical characteristics, but it might also be of fonts that do not exist at present and are in no way related from image to image. These three major issues have to be handled in this research work.
3.2 Image to Image Translation: Npix2Cpix